OpenStudy (anonymous):

how does u*(c-u)=c^2/4 - (c/2-u)^2

OpenStudy (anonymous):

As \[(c/2-u)^2=c^2/4+u^2-cu \implies c^2/4-(c/2-u)^2=-u^2+cu=u(c-u)\] It's just a clever rewrite.